OpenStudy (anonymous):

Solve for x. I will post it equation form in the comments (2)5^(2x-9) -250

OpenStudy (anonymous):

\[(2)5^{2x-9} - 250\]

OpenStudy (anonymous):

Is that - an =, or is this being equated to 0 (essentially equivalent), or something else?

OpenStudy (anonymous):

sorry is an = sign my printer didnt print it off right

OpenStudy (anonymous):

\[(2)5^{2x-9} = 250\]

OpenStudy (anonymous):

Alright. Obvious first step: divide through by 2. Then tell me if you can think of anything that you can do to terms in the form a^x, for example.

OpenStudy (anonymous):

sorry so do you mean 250/2. hmmm, I think it has something to do with using log because thats the opposite of exponents correct?

OpenStudy (anonymous):

I am not exactly sure what you mean by using a^x

OpenStudy (anonymous):

Yeah, I was simply alluding to logs. What base should you take logs in, looking at the left hand side (5 raised to a function of x) and the right hand side (125)?

OpenStudy (anonymous):

i think its 5, but I keep having trouble with this because I dont have a good understanding of using bases. I was reading that my calculator only takes bases of 10

OpenStudy (anonymous):

Yes, it is 5. Your calculator may only calculate logarithms in base 10 and e, but that doesn't matter, since you don't need to use your calculator for this question.

OpenStudy (anonymous):

\[5^{2x - 9} = 125 \to (2x-9)\log_5{5} = \log_5{125}\]By the log relationship:\[\log{a^b} = b \log{a}\]

OpenStudy (anonymous):

so would I come to the equation (2x-9)5 = 5^3?

OpenStudy (anonymous):

You can't simply get rid of the logs. I have taken the log (base 5) of both sides. By the relationship I noted, the right side would be 3log(5). You should also note that log(5) (where log is in base 5) is 1.

OpenStudy (anonymous):

but how would you break it down to solve for x?

OpenStudy (anonymous):

Noting that log base 5 of 5 is 1, the version of the equation I posted simplifies to:\[(2x - 9) = 3\]I imagine you can solve this by yourself. I have to leave now, so I can't provide further elaboration on the working of logs, but if you still don't get it I'm sure somebody else can fill you in.

OpenStudy (anonymous):

okay thank you!